OpenStudy (anonymous):

If f(x) = −x − 1 and g(x) = 4x − 1, find f[g(−2)] −10 −9 −13 8

OpenStudy (anonymous):

g(-2) = 4*-2 - 1 = -9 f[g(−2)] = f(-9) = 9 - 1 = 8

OpenStudy (anonymous):

thanx